OpenStudy (anonymous):

Let f(x)=5x-4 and g(x)=x^2-1. Find each value. 1) \[f(g(\sqrt{6}))\] 2) f(g(0)) 3) \[g(f(4/5))\]

OpenStudy (anonymous):

@dumbcow can you help

OpenStudy (anonymous):

how 'bout me?

OpenStudy (anonymous):

you can too :)

OpenStudy (anonymous):

f(x)=5x-4 and g(x)=x^2-1 #1 f(g(sqrt6)) = lets find g(sqrt6). Can you calculate it?

OpenStudy (dumbcow):

dpalnc will help you :)

OpenStudy (anonymous):

(square root of 6)^2-1=5?

OpenStudy (anonymous):

good. now with this answer, stick it into f(x)

OpenStudy (anonymous):

21

OpenStudy (anonymous):

ok. #2 looks easy what's your answer?

OpenStudy (anonymous):

take your time...

OpenStudy (anonymous):

I got -9

OpenStudy (anonymous):

good. the third one shouldn't be too tough.. what do you have for f(4/5)?

OpenStudy (anonymous):

-1

OpenStudy (anonymous):

is that your answer for #3 or my previous question?

OpenStudy (anonymous):

#3

OpenStudy (anonymous):

then you're done! good job. remember in composition of functions, figure out the "inside" function first then put it on the "outside" function.